OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 4x2 - 32x + 16 = 0

hero (hero):

4x^2 - 32x + 16 = 0 4(x^2 - 8x + 4) = 0 x^2 - 8x + 4 = 0 x^2 - 8x = -4 x^2 - 8x + 16 = 16 - 4 (x - 4)^2 = 12

hero (hero):

To get the 16... (b/2)^2 = (-8/2)^2 = (-4)^2 = 16

OpenStudy (ybarrap):

$$ \tt \bf\\ \Large4x2 - 32x + 16 = 0\\ \Large4(x^2 - 8x + 4) = 0\\ \Large4(x^2 + bx + c) = 0\\ \Large\Large\text{Use } \Large(x+b/2)^2+c-(b/2)^2=0~\text{,with b = -8 and c = 4}\\ \Large\text{So, }(x+b/2)^2=(b/2)^2-c $$